OpenStudy (rebeccapink6398):

What is the perimeter of a rectangle with a length of 7 3/4 inches and a width of 4 1/3 inches? @Nikia_perez

OpenStudy (anonymous):

your answer is 22.

OpenStudy (rebeccapink6398):

they don't have 22

OpenStudy (anonymous):

p=2(l+w)

OpenStudy (rebeccapink6398):

I know that's wat I thought...

OpenStudy (anonymous):

2( 7 3/4 +4 1/3 )=145/6=24.1667

OpenStudy (anonymous):

what are your choices?

OpenStudy (rebeccapink6398):

A. 24 1/6 B.24 1/3 C.11 4/7 D.12 1/12

OpenStudy (anonymous):

Then your answer is A

OpenStudy (rebeccapink6398):

I thought so lol

OpenStudy (anonymous):

145/6 = 24+1/6

OpenStudy (rebeccapink6398):

@robtobey thhx thx thx thx so much now I know how to do it!!! :D